What is the coupon rate of a

ten-year,

$10,000

bond with semiannual coupons and a price of

$8,135.13​,

if it has a yield to maturity of

7.57.5​%?
Answer #1

Number of periods = 10 * 2 = 20

Rate = 7.5% / 2 = 3.75%

Price = Coupon * [1 - 1 / (1 + r)n] / r+ FV / (1 + r)n

8,135.13 = Coupon * [1 - 1 / (1 + 0.0375)20] / 0.0375+ 10000 / (1 + 0.0375)20

8,135.13 = Coupon * [1 - 0.478892] / 0.0375 + 4,788.923421

8,135.13 = Coupon * 13.896213 + 4,788.923421

Coupon = 240.80

Annual coupon = 240.80 * 2 = 481.6

Coupon rate = (481.6 / 10000) * 100

Coupon rate = 4.816%